Session: Oct 2016 (session 3)
Wavelength: 18 cm


To give feedback on this experiment, click here

Experiment Feedback

Badary (Sorokovikov, Slepkova, Lysakova, Zhiritskaya at Tue Nov 01 17:12:27 2016) Success:

Tianma65 (Bo Xia at Tue Nov 01 18:02:45 2016) Success:
EL limited during scan no0256 & no0258.

Medicina (orfei at Tue Nov 01 19:01:13 2016) Minor failures:
Tsys not available. Some scans lost due to repeated crashes of FS.

Irbene (Arturs ORbidans at Tue Nov 01 19:12:13 2016) Success:
antenna not on source at the exact beginning of some scans

Zelenchukskaya (Shatilov A. at Tue Nov 01 19:37:16 2016) Success:

Svetloe (Isaenko O. at Tue Nov 01 19:37:48 2016) Minor failures: ANT
Scans No0017-No0063 failed due to antenna problem.

Hartebeesthoek (Jonathan Quick at Tue Nov 01 19:53:02 2016) Success:
Secondary reflector mispositioned slightly for first 7 scans. No other known problems except for the usual RFI. Several scans were recorded after the source had already set.

Westerbork (Geert Kuper at Wed Nov 02 09:57:53 2016) Success:

Effelsberg (Uwe at Wed Nov 02 11:50:27 2016) Success:

Onsala (Jun Yang at Thu Nov 03 12:00:48 2016) Success:
Due to strong winds, the 25m telescope was stowed during 9:48 -- 16:39. There was a fmout-gps jump triggered by missing two PPS pulses and also reported at 9:53 UT by FS command setcl. The problem with PPS was fixed at 10:20 UT by running FMSET to resync the whole VLBI system.

Torun (Pawel Wolak at Mon Nov 07 09:17:12 2016) Success: